You purchase 75 shares of stock at $ 75 per share. The next day the stock drops 12%; however, a week later it increases in value by 5%. If you immediately sell your 75 shares, how much will you lose or gain (a) in dollar amount and (b) in percent?


1
Expert's answer
2021-06-07T03:31:11-0400

Total number of shares = 75

Price per share = $75

Total amount "= 75 \times 75 = \$5625"

The next day the stock drops 12%

Price per share "= 75 \times (1-0.12) = 75 \times 0.88 = \$66"

However, a week later it increases in value by 5%.

Price per share = "66 \times (1+0.05) =\$69.3"

If you immediately sell your 75 shares:

Total amount for selling "=75 \times 69.3 =\$5197.5"

Difference between final and initial amount = 5197.5 -5625 = -427.5

(a) You will lose $427.5

(b) Percent change "= \frac{427.5}{5625}\times 100 = 7.6 \; \%"